OpenStudy (anonymous):

An orchard operator must dilut 11 quarts of a 60% insecticide solution by adding water. How many quarts of water should be added to get a mixture that is 2% insecticide?

OpenStudy (amistre64):

11(.60-.02)/(1-.60) i think ... might wanna dbl check it tho

OpenStudy (amistre64):

might be 1-.02 tho

OpenStudy (anonymous):

the answer is suppose to be 319 quarts but I cant figure it out

OpenStudy (amistre64):

w + 11(.6) = (w+11)(.02) w+6.6 = w.02 + 2.2 4.4 = w(.02 - 1) 4.4/(.98) = w maybe?

OpenStudy (anonymous):

i dont think thats how

OpenStudy (amistre64):

it aint, i messed it up someplace :)

OpenStudy (amistre64):

w+11 = total quarts w(1) + 11(.60) = (w+11)(.02) should be what they are asking right?

OpenStudy (amistre64):

should the water have a concentration of 1 or 0?

OpenStudy (anonymous):

it does not say, it only gives me the answere 319 quarts

OpenStudy (amistre64):

319 if water is 0 concentration

OpenStudy (amistre64):

w+11 = total quarts w(0) + 11(.60) = (w+11)(.02) 6.6 = w.02 + .22 6.6 - .22 = w.02 w = 6.38/.02 = 638/2 = 319